OpenStudy (anonymous):

how do you solve x^4-4x^2+3=0

OpenStudy (anonymous):

Factor: (x^2 - 1)(x^2 - 3) = 0

OpenStudy (anonymous):

Now, set each factor equal to 0, and find all 4 values of x.

OpenStudy (anonymous):

All four? Why wouldn't it be two?

OpenStudy (anonymous):

x^2 - 1 = 0 (x-1)(x+1) = 0 x = 1, x = -1 x^2 - 3 = 0 x = sqrt(3) and -sqrt(3) Thats why I made sure to mention "4 zeros", as it may be overlooked.

OpenStudy (anonymous):

x^2 - 3 = 0 x^2 = 3 x = sqrt(3) and - sqrt(3)

OpenStudy (anonymous):

Alright. That makes more sense. Thanks

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

 @Kate2997